The length of a room is 4 feet less than twice its width. The perimeter of the room is 58 feet. What is the length of the room? How Would I solve this ?
Answer by checkley79(3341)

L=2W-4
PERIMETER=2L+2W
58=2(2W-4)+2W
58=4W-8+2W
58=6W-8
6W=58+8
6W=66
W=66/6
W=11 ANS. FOR THE WIDTH.
L=2*11-4
L=22-4
L=18 ANS. FOR THE LENGTH.
PROOF:
58=2*18=2*11
58=36+22
58=58